OpenStudy (anonymous):

A swimmer who achieves a speed of 0.80 m/s in still water swims directly across a river 86m wide. the swimmer lands on the far shore at a position 54m downstream from the starting point. Determine the direction the swimmer would have to aim to land directly across from the starting position?

OpenStudy (anonymous):

So here is what I think the diagram looks like... So to get directly across the lake the angle would be \[\tan \theta = opposite/adjacent\] so \[ \theta\]= arctangent (86/54).. = 57.875 degrees so the other angle is 180 - (57.875 + 90) = 32.125 degrees. So I guess the swimmer has to aim 32 degrees to his right. I think that might be the answer because there is no current or anything....